# Enhancing variational quantum algorithms by balancing training on classical and quantum hardware

### v1

@@ Abstract
Quantum computers offer a promising route to tackling problems that are classically intractable such as in prime-factorization, solving large-scale linear algebra and simulating complex quantum systems, but require fault-tolerant quantum hardware. On the other hand, variational quantum algorithms (VQAs) have the potential to provide a near-term route to quantum utility or advantage, and is usually constructed by using parametrized quantum circuits (PQCs) in combination with a classical optimizer for training. Although VQAs have been proposed for a multitude of tasks such as ground-state estimation, combinatorial optimization and unitary compilation, there remain major challenges in its trainability and resource costs on quantum hardware. Here we address these challenges by adopting Hardware Efficient and dynamical LIe algebra Supported Ansatz (HELIA), and propose two training schemes that combine an existing $\mathfrak{g}$-sim method (that uses the underlying group structure of the operators) and the Parameter-Shift Rule (PSR). Our improvement comes from distributing the resources required for gradient estimation and training to both classical and quantum hardware. We numerically test our proposal for ground-state estimation using Variational Quantum Eigensolver (VQE) and classification of quantum phases using quantum neural networks. Our methods show better accuracy and success of trials, and also need fewer calls to the quantum hardware on an average than using only PSR (upto 60% reduction), that runs exclusively on quantum hardware. We also numerically demonstrate the capability of HELIA in mitigating barren plateaus, paving the way for training large-scale quantum models.
@@ I Introduction
Quantum computing holds great promise for tackling problems that are intractable for classical computers or would take them years to solve, such as simulating natural systems [1], prime factorization [2], solving linear equations [3], machine learning tasks [4, 5, 6], optimization [7] and quantum chemistry [8, 9]. Despite recent breakthroughs in implementing quantum error correction [10], it may still take many years [11, 12, 13] to develop fault-tolerant quantum hardware. Current quantum devices face significant challenges, including number of qubits, qubit coherence time and gates with limited fidelity [10, 14, 15, 16, 17, 18, 19].
@@
Limitations of current quantum computing hardware has spurred focus towards variational quantum algorithms (VQA). These are hybrid quantum-classical approaches that utilize parameterized quantum circuits (PQC) to address challenges across diverse fields, including machine learning, optimization, and ground-state energy calculations. VQAs are especially suited for noisy intermediate-scale quantum (NISQ) devices, which have a limited qubit count and noisy (unprotected) operations. The core idea is to train quantum circuits with tunable parameters and optimizing these parameters using a classical optimizer for an objective function related to the problem being solved.
@@
In the training process of VQAs, the parameters of PQC are updated via either gradient-free or gradient-based optimization techniques. The gradient-free methods [20, 21, 22, 23, 24, 25] are preferred when calculation of gradient is challenging or the loss function is not differentiable, but their performance might not scale well with problem size and system noise [26, 27]. In contrast, the literature on gradient-based training of quantum machine learning (QML) primarily emphasizes the efficient computation of gradients of parameters of trainable quantum circuits on quantum hardware. The most commonly employed techniques include the parameter-shift rule (PSR) [28] and its various generalization or extensions [29, 30]. PSR requires running two or more quantum circuit evaluations with shifted parameter values, for each trainable parameter in the circuit, and it run exclusively on quantum hardware.
Classical simulation methods can be an effective alternative for advancing quantum technologies by emulating quantum systems using classical resources. Goh et al. proposed an efficient circuit simulation and gradient computation method called $\mathfrak{g}$-sim , reducing the resource burden leveraging the Lie-algebraic structure of variational quantum circuits [31]. Clifford Perturbation theory has been recently put forward to approximate operator expectation values in near-Clifford circuits [32]. These techniques are crucial for adapting classical backpropagation methods in a quantum context [33].
@@
Optimization in quantum computing is often challenging due to several factors. A major hurdle being the barren plateau (BP) phenomenon [34], where gradients vanish exponentially as qubit count increases  , making it difficult to find numerically meaningful updates for large circuits. Additionally, quantum measurements are inherently stochastic requiring many runs to obtain accurate estimates and thus potentially slowing down optimization. Several BP-free models have been proposed in literature like quantum convolutional neural networks [35, 36] and permutation-equivariant quantum neural networks [37]. Recent research also points towards the possibility that all BP-free quantum models might indeed be efficiently classically simulable [38], raising doubts about achieving any quantum advantage in the future.
Classical algorithms for simulating general quantum circuits do exist, but they suffer from exponential computational and memory overhead [39]. Over the years, several methods have been developed for accurate and efficient classical simulation of certain quantum circuits. These methods often approximate the circuits in specific regimes or restrict the types of gates and measurements allowed, leveraging special structures in the problems. These include MPS-based tensor networks methods that can simulate shallow noisy circuits for hundreds of qubits with limited gate fidelity [40]. $\mathfrak{g}$-sim is another method that offers an efficient simulation algorithm. This method is effective when both the generators of the ansatz and the measurement operator are within a dynamical Lie algebra (DLA) [31], whose dimension scales polynomially with the number of qubits. We will explain more about the $\mathfrak{g}$-sim method in a later section as it is central to our ansatz construction.
Recent advancements in classical simulation methods have significantly enhanced our ability to estimate and optimize quantum circuits, even in the presence of noise. Fontana et.al proposed low weight efficient simulation algorithm (LOWESA) to classically estimate expectation values of parametrized quantum circuits in noisy hardware [41]. The authors recreate a classical surrogate of cost landscape in the presence of device noise by ignoring Pauli terms beyond a certain frequency threshold. Rudolph et.al extended this to noiseless quantum simulation and classically simulated the 127-qubit quantum utility experiment [42, 43] while Angrisani et. al further demonstrated provable guarantees for most noiseless quantum circuits [44]. Recently, Lerch et.al showed it is always possible to identify patches for which similar classical surrogate can be generated [45]. These techniques should be accounted for when constructing variational models with potential for quantum utility.
@@
The optimization landscape in quantum circuits is also highly non-convex, with numerous local minima and saddle points that can trap algorithms and complicate the search for global minima [46, 47, 48, 49]. Additionally, quantum measurements are inherently stochastic requiring many runs to obtain accurate estimates and thus potentially slowing down optimization. Even for models that are trainable, computing gradients for quantum circuits is often quantum resource-intensive as methods such as the parameter-shift rule or finite differences necessitate multiple circuit executions, which scale linearly [33] with the number of parameters and shots. These factors, coupled with hardware limitations such as qubit connectivity and decoherence [10, 14, 15, 16, 17, 18, 19], make effective gradient-based optimization in quantum computing a complex task, frequently driving the need for alternative or hybrid methods.
To mitigate the high quantum resource costs associated with training VQAs, we propose a hybrid method that integrates $\mathfrak{g}$-sim and the PSR. The gradient estimation task is delegated to both classical and quantum hardware. This approach aims to reduce the number of Quantum Processing Unit (QPU) calls, thereby saving resources. Our research focuses on training quantum models, emphasizing the use of specific ansätze and customized training schemes to enhance the training process and improve accuracy in tasks like ground state estimation using VQE and classification using quantum neural networks (QNNs). While this study does not evaluate the practical utility of these models or their performance relative to classical counterparts, it aims to explore the potential of scalable and accurate trainable quantum models. Our hope is that with increase in the size and accuracy of the trainable models we can have enough overlap with useful quantum models and move towards quantum utility.
@@
We will expand on the main results in the coming sections. The paper is divided as follows. We introduce the relevant background literature in Sec. II. We then discuss our contributions in terms of the choice of quantum circuits and two training schemes namely Alternate and Simultaneous in Sec. III. In Sec. IV we demonstrate the improvement in various metrics through extensive numerics. Finally, we end with the Conclusion and Outlook in Sec. V.
@@ II.1 Variational Quantum Algorithms
VQAs involves an $n$-qubit state $\rho$ (usually encoding problem data), acting on a Hilbert space $\mathcal{H}={(\mathbb{C}^{2})}^{\bigotimes n},$ which undergoes evolution through a parameterized quantum unitary. The unitary transformation is expressed as a sequence of $L$ parameterised unitary gates
$U(\boldsymbol{\theta})=\prod_{i=1}^{L}U_{i}(\theta_{i})W_{i}$ (1)
where $\boldsymbol{\theta}=[\theta_{1},\theta_{2},\ldots,\theta_{L}]$ denotes a set of trainable real-valued parameters and $W_{i}$ are non-trainable gates. A slightly more general scenario can be when each $U_{i}(\theta_{i})$ is composed of multiple trainable parametrized unitaries itself. However, we ignore that for ease of notation, and the same ideas can be extended to the general unitaries as well.
@@
$l_{\boldsymbol{\theta}}(\rho,\hat{O})=Tr[U^{\dagger}(\boldsymbol{\theta)}\rho U(\boldsymbol{\theta})\hat{O}].$ (2)
@@
where we minimize some loss function $F$ of the expectation value depending on the task. For example, when finding ground state of a Hamiltonian $H$ using VQE, the loss function is the expectation value of $H$ on the trainable quantum circuit. For classification problem, the loss function can be the mean squared error (MSE) between predicted and true labels. The algorithm iteratively optimizes a cost function using the output of a quantum circuit, enabling the exploration of complex solution spaces that are potentially intractable for classical approaches.
However, despite their promise, VQA face several critical challenges, including barren plateaus in the optimization landscape, limited scalability with increasing problem size, and the detrimental effects of noise and decoherence on circuit fidelity [34, 50, 56]. Overcoming these obstacles necessitates advancements in algorithmic design, optimization techniques, and the development of robust quantum hardware.
Algorithms considered in this work
In this study, we focus on two applications of VQAs, illustrating their potential to address complex quantum problems. The first application investigates the Variational Quantum Eigensolver (VQE), a hybrid quantum-classical algorithm widely used for determining the ground state of quantum systems and calculating their corresponding energies [57]. VQE has become a cornerstone in quantum chemistry and materials science, enabling the efficient simulation of molecular systems and facilitating the discovery of new materials. Recent research has also incorporated VQE into Quantum Error Detection pipeline [58, 59].
The second application delves into quantum phase classification, where VQAs are utilized to identify and distinguish different quantum phases of matter [35, 60, 61]. This application leverages the unique capabilities of quantum circuits to encode and process quantum states, enabling the precise detection of phase transitions and the systematic classification of quantum phases.
II.2 Existing PQCs
@@
×L
Figure 1: Quantum circuit block for YZ linear ansatz. The parametrized $R_{y}$ and $R_{z}$ rotations are applied on each qubit followed by $CNOT$ between neighboring qubits in a linear fashion. The blocks are often repeated to increase entanglement and expressivity
The available options for PQC is quite vast and it is not a priori clear which circuit structure is ideal for the task at hand. We will consider three particular ansatz choices and motivate their use for our ansatz construction.
One often considered ansatz choice is made based on qubit connectivity, to reduce circuit depth, by using the class of Hardware Efficient Ansatz (HEA) [62]. Blocks of single-qubit rotations followed by entangling two-qubit gates are repeated for a chosen number of layers. An example within this class is $YZ$ linear ansatz shown in Fig.1, where each block consists of a $Y$-rotation and a $Z$-rotation applied on each qubit, followed by CNOT gates between neighbouring qubits in a linear fashion.
Commonly selected problem-motivated ansatz include the Unitary Coupled Cluster (UCC) [63], which is widely utilized for investigating ground states of fermionic molecular Hamiltonians, and the Quantum Alternating Operator Ansatz (QAOA) [64, 65, 66], often applied to combinatorial optimization problems.
@@
$U(\boldsymbol{\theta})=\prod_{k}e^{-i\theta_{k}H_{k}}.$ (4)
@@
Building on the insights from [31], this approach can be extended to include terms from the DLA of $\{H_{k}\}_{H_{k}\in H}$, provided the DLA has a dimension that scales polynomially with the number of qubits.
In our examples, we will combine the HVA ansatz with the HEA ansatz. For a detailed overview of the different ansatz options explored in the literature, refer [50].
After selecting an appropriate ansatz based on the architecture and the task, the subsequent step involves determining the optimal parameters of the ansatz to address the problem being solved. This process, referred to as the training method, is elaborated upon in the following section.
II.3 Training Methods
@@
The training process aims to identify the optimal parameters of the selected ansatz for a given architecture. This is an iterative procedure, with each iteration comprising a loss function evaluation followed by a parameter update step. Although several gradient-free methods exist for parameter updation [20, 21, 22, 23], our primary focus here is on gradient-based techniques which require efficient and accurate algorithms to compute the partial derivatives of the cost function for effective training of variational quantum models.

### v2

@@ Abstract
Quantum computers offer a promising route to tackling problems that are classically intractable such as in prime-factorization, solving large-scale linear algebra and simulating complex quantum systems, but potentially require fault-tolerant quantum hardware. On the other hand, variational quantum algorithms (VQAs) are a promising approach for leveraging near-term quantum computers to solve complex problems. However, there remain major challenges in their trainability and resource costs on quantum hardware. Here we address these challenges by adopting Hardware Efficient and dynamical LIe algebra supported Ansatz (HELIA), and propose two training methods that combine an existing classical-enhanced $\mathfrak{g}$-sim method and the quantum-based Parameter-Shift Rule (PSR). Our improvement comes from distributing the resources required for gradient estimation and training to both classical and quantum hardware. We numerically evaluate our approach for ground-state estimation of 6 to 18-qubit Hamiltonians using the Variational Quantum Eigensolver (VQE) and quantum phase classification for up to 12-qubit Hamiltonians using quantum neural networks. For VQE, our method achieves higher accuracy and success rates, with an average reduction in quantum hardware calls of up to 60% compared to purely quantum-based PSR. For classification, we observe test accuracy improvements of up to 2.8%. We also numerically demonstrate the capability of HELIA in mitigating barren plateaus, paving the way for training large-scale quantum models.
@@ I Introduction
Quantum computing holds great promise for tackling problems that are intractable for classical computers or would take them years to solve, such as simulating natural systems [1], prime factorization [2], solving linear equations [3], machine learning tasks [4, 5, 6], optimization [7] and quantum chemistry [8, 9]. Despite recent breakthroughs in implementing quantum error correction [10], it may still take many years [11, 12, 13] to develop fault-tolerant quantum hardware. Current quantum devices face significant challenges, including low number of qubits, qubit coherence time and gates with limited fidelity [10, 14, 15, 16, 17, 18, 19].
@@
Limitations of current quantum computing hardware have spurred focus towards variational quantum algorithms (VQA). These are hybrid quantum-classical approaches that utilize parameterized quantum circuits (PQC) to address challenges across diverse fields, including machine learning, optimization, and ground-state energy calculations. VQAs are especially suited for noisy intermediate-scale quantum (NISQ) devices, which have a limited qubit count and noisy (unprotected) operations. The core idea is to train quantum circuits with tunable parameters and optimizing these parameters using a classical optimizer for an objective function related to the problem being solved.
@@
In VQAs, the parameters of PQC are updated via either gradient-free or gradient-based optimization techniques. The gradient-free methods [20, 21, 22, 23, 24, 25] are preferred when calculation of gradient is challenging or the loss function is not differentiable, but their performance might not scale well with problem size and system noise [26, 27]. In contrast, the literature on gradient-based training of quantum machine learning (QML) primarily emphasizes the efficient computation of gradients of parameters of trainable quantum circuits on quantum hardware. The commonly employed techniques include the parameter-shift rule (PSR) [28] and its various generalization or extensions [29, 30]. PSR requires running two or more quantum circuit evaluations with shifted parameter values for each trainable parameter in the circuit, and it runs exclusively on quantum hardware.
@@
VQAs are often challenging due to several factors. A major hurdle being the barren plateau (BP) phenomenon [31], where gradients vanish exponentially as qubit count increases, making it difficult to find numerically meaningful updates for large circuits. Several BP-free models have been proposed in literature like quantum convolutional neural networks [32, 33] and permutation-equivariant quantum neural networks [34]. Recent research also points towards the possibility that BP-free quantum models might be efficiently classically simulable [35]. These issues need to be taken into consideration when designing new VQA, in order to avoid such pitfalls.
@@
The optimization landscape in quantum circuits is also highly non-convex, with numerous local minima and saddle points that can trap algorithms and complicate the search for global minima [36, 37, 38, 39]. Additionally, quantum measurements are inherently stochastic requiring many runs to obtain accurate estimates and thus potentially slowing down optimization. Even for models that are trainable, computing gradients for quantum circuits is often quantum resource-intensive as methods such as the PSR or finite differences necessitate multiple circuit executions, which scale linearly [40] with the number of parameters and shots. These factors, coupled with hardware limitations such as qubit connectivity and decoherence [10, 14, 15, 16, 17, 18, 19], make effective gradient-based optimization in quantum computing a complex task, frequently driving the need for alternative or hybrid methods.
Classical algorithms for simulating general quantum circuits do exist, but they suffer from exponential computational and memory overhead [41]. Several methods have been developed for efficient classical simulation of certain classes of quantum circuits, leveraging special structures in the problems. These include Matrix Product State-based tensor networks methods that can simulate shallow noisy circuits for hundreds of qubits with limited gate fidelity [42]. Clifford Perturbation theory has been recently put forward to approximate operator expectation values in near-Clifford circuits [43]. $\mathfrak{g}$-sim is another method that offers an efficient simulation algorithm, based on the study of the Lie group and the associated Lie algebra $\mathfrak{g}$, which is generated by the PQC. This method is effective when both the generators of the ansatz and the measurement operator are within a dynamical Lie algebra (DLA) [44], whose dimension scales polynomially with the number of qubits. Initially porposed by Somma et al. [45], the techniques were reframed in a modern context tailored to the quantum computing community by Goh et al. [44]. We will explain more about the $\mathfrak{g}$-sim method in a later section as it is central to our ansatz construction.
Although simulating arbitrary quantum circuits is hard, recent advancements in classical simulation methods have significantly enhanced our ability to estimate quantum circuits in specific parameter regions and in the presence of noise. Fontana et al. proposed low weight efficient simulation algorithm (LOWESA) to classically estimate expectation values of PQCs in noisy hardware [46]. The authors recreate a classical surrogate of the cost landscape in the presence of device noise by ignoring Pauli terms beyond a certain frequency threshold. Rudolph et al. extended this to noiseless quantum simulation and classically simulated the 127-qubit quantum utility experiment [47, 48] while Angrisani et. al further demonstrated provable guarantees for most noiseless quantum circuits [49]. Recently, Lerch et al. showed it is always possible to identify patches for which similar classical surrogate can be generated [50]. These techniques should be accounted for when constructing variational models with potential for quantum utility.
Figure 1: Overview of our main contributions (a) We propose Hardware Efficient and dynamical LIe algebra supported Ansatz (HELIA) composed of two blocks of gates $U_{q}$ and $U_{\mathfrak{g}}$ whose gradients can be obtained using parameter-shift rule and $\mathfrak{g}$-sim respectively. The resources required for gradient evaluation of the full ansatz can hence be delegated to both quantum and classical hardware. We further propose two training methods: Alternate and Simultaneous which benefit from the hybrid gradient of HELIA. (b) To evaluate the gradients of $U_{\mathfrak{g}}$, we use a $\mathfrak{g}$-sim based method. The operators of a dynamical Lie algebra are measured after applying $U_{q}$ (using QPU in the leftmost block), which is then passed on to classical hardware that evaluates the cost function and gradients using $\mathfrak{g}$-sim (using CPU in the middle block). Further details of our contribution are elaborated in Sec. III. Using our proposed ansatz and training methods we are able to reduce quantum hardware usage, improved accuracy and mitigate Barren Plateaus as elaborated in Sec. IV.
We focus on enhancing quantum model training through tailored ansätze and customized training methods, improving accuracy in tasks such as ground-state estimation with the VQE and quantum phase classification using quantum neural networks (QNNs). To mitigate the high quantum resource costs associated with training VQAs, we propose Hardware Efficient and dynamical LIe algebra supported Ansatz (HELIA) and two hybrid methodsAlternate and Simultaneousthat integrate $\mathfrak{g}$-sim and PSR (Fig. 1). Our gradient estimation task is delegated to both classical and quantum hardware to reduce the number of Quantum Processing Unit (QPU) calls, thereby saving resources.
While this study does not evaluate the practical utility of these models or their performance relative to classical counterparts, it aims to explore the potential of scalable and accurate trainable quantum models. Our findings suggest that increasing the scale and accuracy of trainable models could bridge the gap toward practical quantum utility.
@@
The remaining sections are divided as follows. We introduce the relevant background literature in Sec. II. We then discuss our contributions in terms of the choice of quantum circuits and two training methods namely Alternate and Simultaneous in Sec. III. In Sec. IV we demonstrate the improvement in various metrics through extensive numerical simulations. Finally, we end with the Conclusion and Outlook in Sec. V.
@@ II.1 Variational Quantum Algorithms
VQAs involve an $n$-qubit state $\rho$ (usually encoding problem data), acting on a Hilbert space $\mathcal{H}={(\mathbb{C}^{2})}^{\bigotimes n},$ which undergoes evolution through a parameterized quantum unitary. The unitary transformation is expressed as a sequence of $L$ parameterised unitary gates
$U(\boldsymbol{\theta})=\prod_{i=1}^{L}U_{i}(\boldsymbol{\theta}_{i})W_{i}$ (1)
where $\boldsymbol{\theta}=[\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2},\ldots,\boldsymbol{\theta}_{L}]$ denotes a set of trainable real-valued parameters and $W_{i}$ are non-trainable gates. An example of such case is shown in Fig. 2, where $U_{i}(\boldsymbol{\theta}_{i})$ is be composed of multiple trainable single-qubit gates, while the non-trainable block contains entangling CNOT gates.
@@
[…]
@@
where we minimize some loss function $F$ of the expectation value depending on the task.
The form of the loss function in Eq. (3) can vary depending on the task. For example, when estimating ground states the loss function can be the expectation value of a Hamiltonian, and finding the ground state is equivalent to minimizing this expectation value. For classification task, the loss function can be mean squared error between predicted labels from a quantum model and the true labels obtained from a dataset. In both cases, the algorithm iteratively optimizes a cost function using the output of a quantum circuit, enabling the exploration of complex solution spaces that are potentially intractable for classical approaches.
Despite their promise, VQA face several critical challenges, including barren plateaus in the optimization landscape, limited scalability with increasing problem size, and the detrimental effects of noise and decoherence on circuit fidelity [31, 51, 57]. Overcoming these obstacles necessitates advancements in algorithmic design, optimization techniques, and the development of robust quantum hardware.
@@
[…]
Figure 2: Quantum circuit block for YZ linear (HEA) ansatz. The parametrized $R_{y}$ and $R_{z}$ rotations are applied on each qubit followed by $CNOT$ between neighboring qubits in a linear fashion. The blocks are often repeated to increase entanglement and expressivity.
The available options for PQC in VQAs is quite vast and it is not a priori clear which circuit structure is ideal for the task at hand. One common choice is Hardware Efficient Ansatz (HEA), which consists of blocks of single-qubit rotations followed by entangling two-qubit gates repeated for a chosen number of layers [58]. An example within this class is the $YZ$ linear ansatz shown in Fig. 2, where each block consists of a $Y$-rotation and a $Z$-rotation applied on each qubit, followed by CNOT gates between neighbouring qubits in a linear fashion. Commonly selected problem-motivated ansatz include the Unitary Coupled Cluster (UCC) [59], which is widely utilized for investigating ground states of fermionic molecular Hamiltonians, and the Quantum Alternating Operator Ansatz (QAOA) [60, 61, 62], often applied to combinatorial optimization problems.
@@
$U(\boldsymbol{\theta})=\prod_{i}e^{-i\theta_{i}H_{i}},$ (4)
@@
and can repeated for a chosen number of layers. Building on the insights from Ref. [44], this approach can be extended to include terms from the dynamical Lie Algebra (DLA) of $\{iH_{i}\}_{H_{i}\in H}$, provided the DLA has a dimension that scales polynomially with the number of qubits. We elaborate on the details of DLA in Sec. II.2.2. In this study, we will combine the HVA ansatz with the HEA ansatz. For a detailed overview of the ansatz options explored in the literature, we refer to Ref. [51].
II.2 Training Methods
@@
After selecting an appropriate ansatz, the subsequent step involves determining the optimal parameters of the ansatz to address the problem being solved. This is an iterative procedure, with each iteration comprising a loss function evaluation followed by a parameter update step. Although several gradient-free methods exist for parameter update [20, 21, 22, 23], here we focus on gradient-based techniques, which require efficient and accurate algorithms to compute the partial derivatives of the cost function.